OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10)and (-4, 0)? Answer - -

OpenStudy (anonymous):

Slope formula between two points

OpenStudy (anonymous):

slope = \[\Delta y / \Delta x\] =(-10-0)/(0-(-4)) =-10/4 =-5/2